# A note on Lie and Jordan structures of Leavitt path algebras

### v1

@@ 1. Introduction
Every associative algebra $\mathcal{A}$ over a field $K$ can give rise to two new structures: the Lie algebra, denoted by $\mathcal{A}^{-}$, and the Jordan algebra, denoted by $\mathcal{A}^{\circ}$ over $K$. These structures can be obtained by defining new products within $\mathcal{A}$, namely the Lie bracket $[x,y]=xy-yx$ for Lie algebras and the Jordan product $x\circ y=xy+yx$ for Jordan algebras respectively, for every $x,y\in\mathcal{A}$. A reasonable question arises: Are the properties of $\mathcal{A}$ as a ring mirrored in its Lie or Jordan algebra counterparts? Specifically, does the simplicity of $\mathcal{A}$ as a ring guarantee its simplicity as a Lie or Jordan algebra? In a seminal paper [9] by Herstein, this question was answered affirmatively. Assume further that $\mathcal{A}$ is equipped with an involution $\star$; that is, a map ${}^{\star}:\mathcal{A}\to\mathcal{A}$ satisfying all below conditions:
@@
[…]
In this paper, we consider the case when $\mathcal{A}$ is the Leavitt path algebra $L_{K}(E)$ of a directed graph $E$ over a field $K$. Within the past few years, the Lie subalgebras of $L_{K}(E)^{-}$ arising from $L_{K}(E)$ have been investigated by many authors. Motivated by the groundbreaking work of Herstein [9], necessary and sufficient conditions for the Lie algebra $[L_{K}(E),L_{K}(E)]$ of $L_{K}(E)^{-}$ to be simple were given in [2] and [3], under the assumption that $E$ is row-finite. In another direction, the paper [12] completely identified the graph $E$ and the field $K$ for the Lie algebra $L_{K}(E)^{-}$ to be solvable.
@@ 2. Preliminaries
Let $E$ be a graph. A vertex $v$ in $E$ is a sink if it emits no edges, while it is an infinite emitter if it emits infinitely many edges. A graph $E$ is said to be row-finite if $E$ contains no infinite emitter. A vertex $v$ is said to be regular if it is neither a sink nor an infinite emitter. A path $\mu$ in a graph $E$ is a finite sequence of edges $\mu=e_{1}e_{2}\dots e_{n}$ with $r(e_{i})=s(e_{i+1})$ for all $11\leq i\leq n-$ 1. We set $s(\mu):=s(e_{1})$ and $r(\mu):=r(e_{n})$. The set of all finite paths in $E$ is denoted by ${\rm Path}(E)$. An exit for a path $\mu=e_{1}\dots e_{n}$ is an edge $e$ such that $s(e)=s(e_{i})$ for some $i$ and $e\neq e_{i}$.
@@ 2.2. Lie and Jordan structures of associative algebras
Every associative algebra $\mathcal{A}$ over a field $K$ is endowed with a Lie algebra structure for the Lie bracket $[a,b]=ab-ba$ for all $a,b\in\mathcal{A}$. We denote this Lie algebra by $\mathcal{A}^{-}$. For $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we defined $[S,T]={\rm span}_{K}\{ab-ba:a\in S,b\in T\}$. A $K$-subspace $\mathcal{L}$ of $\mathcal{A}$ is called a Lie subalgebra of $\mathcal{A}^{-}$ if $[a,b]\in\mathcal{L}$ for all $a,b\in\mathcal{L}$. For the Lie subalgebra $\mathcal{L}$ of $\mathcal{A}^{-}$, we define inductively the two series
@@
The Lie subalgebra $\mathcal{L}$ is called Lie solvable (resp., Lie nilpotent) if there exists $n\geq 0$ (resp., $m\geq 0$) such that $\mathcal{L}^{(n)}=0$ (resp., $\mathcal{L}^{m}=0$). If such the smallest $n$ (resp., $m$) exists, we will say that $\mathcal{L}$ is Lie solvable of index $n$ (resp., Lie nilpotent of index $m$).
Similarly, a Jordan product can be defined on $\mathcal{A}$ by $a\circ b=ab+ba$, for all $a,b\in\mathcal{A}$. If $\mathcal{A}$ is viewed as a Jordan algebra for the product $\circ$, we denote it by $\mathcal{A}^{\circ}$. Note that in the case ${\rm char}(K)=2$, the two structures $\mathcal{A}^{\circ}$ and $\mathcal{A}^{-}$ coincides. A $K$-subspace $\mathcal{J}$ of $\mathcal{A}$ is called a Jordan subalgebra of $\mathcal{A}^{\circ}$ if $a\circ b\in\mathcal{J}$ for all $a,b\in\mathcal{J}$. For two $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we defined
@@ 2.3. Homomorphisms of algebras with involutions
It is obvious that if $(\mathcal{A},^{\star})$ is isomorphic to $(\mathcal{B},^{\natural})$ as $K$-algebras with involution, then the Lie subalgebras $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{K}_{\mathcal{B}}$, with the respect involutions and $\natural$ respectively, are also isomorphic as Lie algebras over $K$. Similarly, the Jordan subalgebras $\mathbf{S}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{B}}$ of $\mathcal{A}^{\circ}$ and $\mathcal{B}^{\circ}$, with the respect involutions and $\natural$ respectively, are also isomorphic as Jordan algebras over $K$.
@@ 3. Lie Solvability and Jordan Solvability of $\mathbb{M}_{n}(\mathcal{A})$
for every matrix $(a_{ij})\in\mathbb{M}_{n}(\mathcal{A})$. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be the Lie subalgebra of $\mathbb{M}_{n}(\mathcal{A})^{-}$ with respect to the involution $\natural$ on $\mathbb{M}_{n}(\mathcal{A})$ as defined in (3.1). In this section, we gives conditions on $n$ and $\mathcal{A}$ for $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ to be Lie solvable. First, we fix the following notation: In $\mathbb{M}_{n}(\mathcal{A})$, we denote by $E_{ij}$ the matrix whose $(i,j)$-entry is $1$ and the other entries are 00. The following proposition is the main result of this section:
@@
Let $K$ be a field, and $\mathcal{A}$ be a unital $K$-algebra with an involution . Assume further that $\mathcal{A}$ is an integral domain. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be as defined above. Then, the following assertions holds:
@@
    Assume that ${\rm char}(K)\neq 2$. If the involution on $\mathcal{A}$ is trivial, then $[\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$; while if the involution on $\mathcal{A}$ is non-trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable.
@@
(3.4) $[X_{1},X_{2}]=\left(\begin{array}[]{cc}s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}&0\\[8.5359pt] 2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)&s_{2}^{\natural}s_{1}-s_{1}^{\natural}s_{2}\end{array}\right),\text{ where }$
@@
(b) and (c): Assume that ${\rm char}(K)=2$. As $02(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)=$ 0, it follows from (3.4) that $[X_{1},X_{2}]$ is a diagonal matrix. Similarly, we also obtain that $[X_{3},X_{4}]$ is also a diagonal matrix. Thus $[[X_{1},X_{2}],[X_{3},X_{4}]]=0$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(3)}=0$. We divide our situation into two possible cases:
Case 1. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is trivial. As $b_{i}^{\natural}=b_{i}$ and $v_{i}^{\natural}=v_{i}$, we get from (3.3) that $r_{i}=0$. It follows from (3.2) that $X_{i}=\left(\begin{array}[]{cc}0&s_{i}\\ -s_{i}^{\natural}&0\end{array}\right)$, and so $[X_{1},X_{2}]=0$; or equivalently, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(2)}=0$. In other words, $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 2. This completes the proof of the second assertion of (c).
Case 2. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is non-trivial. In this case, suitable choices of the matrices $A_{1},A_{2},B_{1},B_{2}$ yield that $[X_{1},X_{2}]\neq 0$. For example, if we choose $a_{1}=c_{1}$, $a_{2}=c_{2}$, $b_{1}=1$, $u_{1}\neq w_{1}$, $w_{1}=w_{2}=0$, $0\neq u_{1}=u_{2}\in K$, $b_{2}\neq b_{2}^{\natural}$, then it follows from (3.5) that $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}=(b_{2}^{\natural}-b_{2})u_{1}^{2}\neq 0$. Therefore, we get from (3.4) that $[X_{1},X_{2}]\neq 0$. It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 3. This completes the proof of (c).
@@
Case 3. ${\rm char}(K)\neq 2$ and the involution on $\mathcal{A}$ is trivial. Because $a^{\natural}=a$, the equation $a=-a^{\natural}$ is equivalent to $02a=$ 0, and hence $a=0$. Similarly, we also obtain that $c=0$. Moreover, as $b^{\natural}=b$, we finally get that
@@
It follows that $[\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$.
@@
    Suppose that $n=2$. If ${\rm char}(K)\neq 2$, then $[\mathbf{K}_{\mathbb{M}_{n}(K)},\mathbf{K}_{\mathbb{M}_{n}(K)}]=0$; and, if ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is Lie solvable of index $2$ but not Lie nilpotent.
@@
Let $K[x,x^{-1}]$ be the Laurent polynomial $K$-algebra. Let $\mathbf{K}_{\mathbb{M}_{n}(K[x,x^{-1}])}$ be the Lie subalgebra of skew-symmetric elements of $\mathbb{M}_{n}(K[x,x^{-1}])^{-}$ with respect to the involution defined by ((((3.7)))). Then, the following assertions hold:
@@ 4. Lie Solvability and Jordan Solvability of $L_{K}(E)$
Throughout of this section, we will denote by $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$with respect to standard involution on $L_{K}(E)$ as defined before, respectively. The main aim of this section is to provide a necessary and sufficient condition on the graph $E$ and the field $K$ for $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ to be Lie and Jordan solvable, respectively. We start this section a lemma which determines a $K$-basis of $L_{K}(E)$:
@@
Let $E$ be a graph and $K$ a field. Put $\mathscr{A}=\{\lambda\nu^{*}\;|\;\lambda,\nu\in{\rm Path}(E)\text{ and }r(\lambda)=r(\nu)\}$. For each $v\in{\rm Reg}(E)$, let $\{e_{1}^{v},\dots,e_{n_{v}}^{v}\}$ be an enumeration of the elements of $s^{-1}(v)$. Then, a $K$-basic of $L_{K}(E)$ is given by the family
@@
The next lemma provides a necessary condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. The ideas used to prove of this lemma come from the proof of [12, Lemma 2.2 and Theorem 2.3]. It can be said that, in $L_{K}(E)$, we can easily choose a suitable subalgebra which is isomorphic to the matrix ring $\mathbb{M}_{3}(K)$. Then, Corollary 3.2 is applied to get all possible cases for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. For convenience of the readers, we still include the proof here.
@@
. Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. Then, $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable if one of the following cases occurs:
@@
[…]
@@
(2) For the proof of this assertion, we assume first that $E$ contains $F_{1}$. Then, $v,f,ef$ are all paths ending at $v$. It is straightforward to check the subalgebra $B$ of $L_{K}(E)$ generated by $v$, $f$, $ef$, $f^{*}$, $f^{*}e^{*}$, $ff^{*}$, $eff^{*}$, $ff^{*}e^{*}$, $eff^{*}e^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map: $eff^{*}e^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $eff^{*}\mapsto E_{12}$, $ff^{*}e^{*}\mapsto E_{21}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$, $ef\mapsto E_{13}$, and $f^{*}e^{*}\mapsto E_{31}$ (see also a similar techniques used in the proof of [1, Lemma 2.6.4]). The restriction of the involution to $B$ is also an involution on $B$ for which $\mathbf{K}_{B}\subseteq\mathbf{K}_{L_{K}(E)}$. Thus, $\mathbf{K}_{B}$ is isomorphic to $\mathbf{K}_{\mathbb{M}_{3}(K)}$ as Lie algebras over $K$. Therefore, by Corollary 3.2, we conclude that $\mathbf{K}_{B}$, and so $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable.
@@
Let $\mathcal{A}$ be a $K$-algebra with an involution , and $\mathbf{J}$ be an indexing set of arbitrary cardinality. Let $I_{j},j\in\mathbf{J}$, be ideals of $\mathcal{A}$ satisfying the property that for each $j$, if $a\in I_{j}$ then $a^{\star}\in I_{j}$. Put $M=\sum_{j\in\mathbf{J}}I_{j}$. If $M=\bigoplus_{j\in\mathbf{J}}I_{j}$, then $\mathbf{K}_{M}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}$. Consequently, $\mathbf{K}_{M}$ is Lie solvable ((((nilpotent)))) if and only if $\mathbf{K}_{I_{j}}$ is Lie solvable ((((nilpotent)))) of a common index for all $j\in\mathbf{J}$.
@@
    $\mathbf{K}_{L_{K}(E_{3})}$ is Lie solvable ((((of index $3$)))) if and only if $\mathrm{char}(K)=2$. Additionally, $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent.

### v2

@@ 1. Introduction
Every associative algebra $\mathcal{A}$ over a field $K$ can give rise to two new structures: the Lie algebra, denoted by $\mathcal{A}^{-}$, and the Jordan algebra, denoted by $\mathcal{A}^{\circ}$ over $K$. These structures can be obtained by defining new products within $\mathcal{A}$, namely the Lie bracket [x,y]=xyyx[x,y]=xy-yx for Lie algebras and the Jordan product $x\circ y=xy+yx$ for Jordan algebras respectively, for every $x,y\in\mathcal{A}$. A reasonable question arises: Are the properties of $\mathcal{A}$ as a ring mirrored in its Lie or Jordan algebra counterparts? Specifically, does the simplicity of $\mathcal{A}$ as a ring guarantee its simplicity as a Lie or Jordan algebra? In a seminal paper [11] by Herstein, this question was answered affirmatively. Assume further that $\mathcal{A}$ is equipped with an involution $\star$; that is, a map $^{\star}:\mathcal{A}\to\mathcal{A}$ satisfying the following conditions:
@@
be the sets of symmetric elements and skew-symmetric elements of $\mathcal{A}$ with respect to $\star$}, respectively. It is obvious that $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{A}}$ are Lie subalgebra and Jordan subalgebra of $\mathcal{A}^{-}$ and $\mathcal{A}^{\circ}$ respectively. These are the most important structures of Lie or Jordan algebras arising from an associative algebra. In particular, when $\mathcal{A}$ is the full matrix ring $\mathbb{M}_{n}(K)$ and is the transpose operation, then $\mathbf{K}_{\mathcal{A}}$ is identified with the Lie algebra of skew-symmetric matrices which corresponds to the Lie group of orthogonal matrices. Under certain assumptions on $\mathcal{A}$, Lie and Jordan structures of $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{A}}$ respectively were intensively investigated by Herstein in a series of papers ([7]-[11]) and by other authors in the references therein.
In this paper, we consider the case when $\mathcal{A}$ is the Leavitt path algebra $L_{K}(E)$ of a directed graph $E$ over a field $K$. The concept of the Leavitt path algebra $L_{K}(E)$ associated with a graph $E$ and coefficients from a field $K$ made its debut in [2], sparking widespread interest among scholars. These algebras were concurrently and separately introduced in [6], despite being published two years following [2]. Within the past few years, the Lie subalgebras of $L_{K}(E)^{-}$ arising from $L_{K}(E)$ have been investigated by many authors. Motivated by the groundbreaking work of Herstein [11], necessary and sufficient conditions for the Lie algebra $L_{K}(E),L_{K}(E)]$ of $L_{K}(E)^{-}$ to be simple were given in [3] and [4], under the assumption that $E$ is row-finite. In another direction, the paper [12] completely identified the graph $E$ and the field $K$ for the Lie algebra $L_{K}(E)^{-}$ to be solvable.
@@ 2. Preliminaries
Let $E$ be a graph. A vertex $v$ in $E$ is a sink if it emits no edges, while it is an infinite emitter if it emits infinitely many edges. A graph $E$ is said to be row-finite if $E$ contains no infinite emitter. A vertex $v$ is said to be regular if it is neither a sink nor an infinite emitter. A path $\mu$ in a graph $E$ is a finite sequence of edges $\mu=e_{1}e_{2}\dots e_{n}$ with $r(e_{i})=s(e_{i+1})$ for all $1\leq i\leq n-1$. We set $s(\mu):=s(e_{1})$ and $r(\mu):=r(e_{n})$. The set of all finite paths in $E$ is denoted by ${\rm Path}(E)$. An exit for a path $\mu=e_{1}\dots e_{n}$ is an edge $e$ such that $s(e)=s(e_{i})$ for some $i$ and $e\neq e_{i}$.
@@ 2.2. Lie and Jordan structures of associative algebras
Every associative algebra $\mathcal{A}$ over a field $K$ is endowed with a Lie algebra structure for the Lie bracket [a,b]=abba[a,b]=ab-ba for all $a,b\in\mathcal{A}$. We denote this Lie algebra by $\mathcal{A}^{-}$. For $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we define $S,T]={\rm span}_{K}\{ab-ba:a\in S,b\in T\}$. A $K$-subspace $\mathcal{L}$ of $\mathcal{A}$ is called a Lie subalgebra of $\mathcal{A}^{-}$ if $a,b]\in\mathcal{L}$ for all $a,b\in\mathcal{L}$. For the Lie subalgebra $\mathcal{L}$ of $\mathcal{A}^{-}$, we define inductively the two series
@@
The Lie subalgebra $\mathcal{L}$ is called Lie solvable (resp., Lie nilpotent) if there exists $n\geq 0$ (resp., $m\geq 0$) such that $\mathcal{L}^{(n)}=0$ (resp., $\mathcal{L}^{m}=0$). If such a smallest $n$ (resp., $m$) exists, we will say that $\mathcal{L}$ is Lie solvable of index $n$ (resp., Lie nilpotent of index $m$).
Similarly, a Jordan product can be defined on $\mathcal{A}$ by $a\circ b=ab+ba$, for all $a,b\in\mathcal{A}$. If $\mathcal{A}$ is viewed as a Jordan algebra for the product $\circ$, we denote it by $\mathcal{A}^{\circ}$. Note that in the case ${\rm char}(K)=2$, the two structures $\mathcal{A}^{\circ}$ and $\mathcal{A}^{-}$ coincide. A $K$-subspace $\mathcal{J}$ of $\mathcal{A}$ is called a Jordan subalgebra of $\mathcal{A}^{\circ}$ if $a\circ b\in\mathcal{J}$ for all $a,b\in\mathcal{J}$. For two $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we define
@@ 2.3. Homomorphisms of algebras with involutions
It is obvious that if $(\mathcal{A},^{\star})$ is isomorphic to $(\mathcal{B},^{\natural})$ as $K$-algebras with involution, then the Lie subalgebras $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{K}_{\mathcal{B}}$, with respect to the involutions and $\natural$ respectively, are also isomorphic as Lie algebras over $K$. Similarly, the Jordan subalgebras $\mathbf{S}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{B}}$ of $\mathcal{A}^{\circ}$ and $\mathcal{B}^{\circ}$, with respect to the involutions and $\natural$ respectively, are also isomorphic as Jordan algebras over $K$.
@@ 3. Lie Solvability and Jordan Solvability of $\mathbb{M}_{n}(\mathcal{A})$
for every matrix $(a_{ij})\in\mathbb{M}_{n}(\mathcal{A})$. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be the Lie subalgebra of $\mathbb{M}_{n}(\mathcal{A})^{-}$ with respect to the involution $\natural$ on $\mathbb{M}_{n}(\mathcal{A})$ as defined in (3.1). In this section, we give conditions on $n$ and $\mathcal{A}$ for $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ to be Lie solvable. First, we fix the following notation: In $\mathbb{M}_{n}(\mathcal{A})$, we denote by $E_{ij}$ the matrix whose $(i,j)$-entry is $1$ and the other entries are 0. The following proposition is the main result of this section:
@@
Let $K$ be a field, and $\mathcal{A}$ be a unital $K$-algebra with an involution . Assume further that $\mathcal{A}$ is an integral domain. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be as defined above. Then, the following assertions hold:
@@
    Assume that ${\rm char}(K)\neq 2$. If the involution on $\mathcal{A}$ is trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$; while if the involution on $\mathcal{A}$ is non-trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable.
@@
(3.4) $X_{1},X_{2}]=\left(\begin{array}[]{cc}s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}&0\\[8.5359pt] 2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)&s_{2}^{\natural}s_{1}-s_{1}^{\natural}s_{2}\end{array}\right),\text{ where }$
@@
(b) and (c): Assume that ${\rm char}(K)=2$. As $2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)=0$, it follows from (3.4) that $X_{1},X_{2}]$ is a diagonal matrix. Similarly, we also obtain that $X_{3},X_{4}]$ is also a diagonal matrix. Thus $X_{1},X_{2}],[X_{3},X_{4}]]=0$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(3)}=0$. We divide our situation into two possible cases:
Case 1. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is trivial. As $b_{i}^{\natural}=b_{i}$ and $v_{i}^{\natural}=v_{i}$, we get from (3.3) that $r_{i}=0$. It follows from (3.2) that $X_{i}=\left(\begin{array}[]{cc}0&s_{i}\\ -s_{i}^{\natural}&0\end{array}\right)$, and so $X_{1},X_{2}]=0$; or equivalently, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(2)}=0$. In other words, $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 2. This completes the proof of the second assertion of (c).
Case 2. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is non-trivial. In this case, suitable choices of the matrices $A_{1},A_{2},B_{1},B_{2}$ yield that $X_{1},X_{2}]\neq 0$. For example, if we choose $a_{1}=c_{1}$, $a_{2}=c_{2}$, $b_{1}=1$, $u_{1}\neq w_{1}$, $w_{1}=w_{2}=0$, $0\neq u_{1}=u_{2}\in K$, $b_{2}\neq b_{2}^{\natural}$, then it follows from (3.5) that $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}=(b_{2}^{\natural}-b_{2})u_{1}^{2}\neq 0$. Therefore, we get from (3.4) that $X_{1},X_{2}]\neq 0$. It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 3. This completes the proof of (c).
@@
Case 3. ${\rm char}(K)\neq 2$ and the involution on $\mathcal{A}$ is trivial. Because $a^{\natural}=a$, the equation $a=-a^{\natural}$ is equivalent to $2a=0$, and hence $a=0$. Similarly, we also obtain that $c=0$. Moreover, as $b^{\natural}=b$, we finally get that
@@
It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$.
@@
    Suppose that $n=2$. If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)},\mathbf{K}_{\mathbb{M}_{n}(K)}]=0$; and, if ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is Lie solvable of index $2$ but not Lie nilpotent.
@@
Let $K[x,x^{-1}]$ be the Laurent polynomial $K$-algebra. Let $\mathbf{K}_{\mathbb{M}_{n}(K[x,x^{-1}])}$ be the Lie subalgebra of skew-symmetric elements of $\mathbb{M}_{n}(K[x,x^{-1}])^{-}$ with respect to the involution defined by ((3.7)). Then, the following assertions hold:
@@ 4. Lie Solvability and Jordan Solvability of $L_{K}(E)$
Throughout this section, we will denote by $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$with respect to standard involution on $L_{K}(E)$ as defined before, respectively. The main aim of this section is to provide a necessary and sufficient condition on the graph $E$ and the field $K$ for $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ to be Lie and Jordan solvable, respectively. We start this section with a lemma which determines a $K$-base of $L_{K}(E)$:
@@
Let $E$ be a graph and $K$ a field. Put $\mathscr{A}=\{\lambda\nu^{*}\;|\;\lambda,\nu\in{\rm Path}(E)\text{ and }r(\lambda)=r(\nu)\}$. For each $v\in{\rm Reg}(E)$, let $\{e_{1}^{v},\dots,e_{n_{v}}^{v}\}$ be an enumeration of the elements of $s^{-1}(v)$. Then, a $K$-basis of $L_{K}(E)$ is given by the family
@@
The next lemma provides a necessary condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. The ideas used to prove this lemma come from the proof of [12, Lemma 2.2 and Theorem 2.3]. It can be said that, in $L_{K}(E)$, we can easily choose a suitable subalgebra which is isomorphic to the matrix ring $\mathbb{M}_{3}(K)$. Then, Corollary 3.2 is applied to get all possible cases for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. For convenience of the readers, we still include the proof here.
@@
Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. Then, $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable if one of the following cases occurs:
@@
    $F_{1}:\bullet\bullet\bulletvveeff$
    $F_{2}:\bullet\bullet\bulletvveeff$
    $F_{3}:\bullet\bulletvveeff$
@@
(2) For the proof of this assertion, we assume first that $E$ contains $F_{1}$. Then, $v,f,ef$ are all paths ending at $v$. It is straightforward to check the subalgebra $B$ of $L_{K}(E)$ generated by $v$, $f$, efef, $f^{*}$, $^{*}e^{*}, ff$$f^{*}$, $^{*}, ffeff^{*}e^{*}$, $eff^{*}e^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map: $eff^{*}e^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $eff^{*}\mapsto E_{12}$, $ff^{*}e^{*}\mapsto E_{21}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$, $ef\mapsto E_{13}$, and $f^{*}e^{*}\mapsto E_{31}$ (see also a similar techniques used in the proof of [1, Lemma 2.6.4]). The restriction of the involution to $B$ is also an involution on $B$ for which $\mathbf{K}_{B}\subseteq\mathbf{K}_{L_{K}(E)}$. Thus, $\mathbf{K}_{B}$ is isomorphic to $\mathbf{K}_{\mathbb{M}_{3}(K)}$ as Lie algebras over $K$. Therefore, by Corollary 3.2, we conclude that $\mathbf{K}_{B}$, and so $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable.
@@
Let $\mathcal{A}$ be a $K$-algebra with an involution $\star$, and $\mathbf{J}$ be an indexing set of arbitrary cardinality. Let $I_{j},j\in\mathbf{J}$, be ideals of $\mathcal{A}$ satisfying the property that for each $j$, if $a\in I_{j}$ then $a^{\star}\in I_{j}$. Put $M=\sum_{j\in\mathbf{J}}I_{j}$. If $M=\bigoplus_{j\in\mathbf{J}}I_{j}$, then $\mathbf{K}_{M}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}$. Consequently, $\mathbf{K}_{M}$ is Lie solvable ((nilpotent)) if and only if $\mathbf{K}_{I_{j}}$ is Lie solvable ((nilpotent)) of a common index for all $j\in\mathbf{J}$.
@@
    $\mathbf{K}_{L_{K}(E_{3})}$ is Lie solvable ((of index 33)) if and only if $\mathrm{char}(K)=2$. Additionally, $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent.